Two angles are complementary if thier sum is90 degrees. If the measure of the firt angle isxdegrees and the measure of the second angle is (3x-2)degrees, find the measure of each angle
Found 3 solutions by vleith, checkley77, Earlsdon:
Answer by vleith(2983)
Answer by checkley77(12844)

x+(3x-2)=90
x+3x-2=90
4x=90+2
4x=92
x=92/4
x=23 for the smaller angle.
3*23-2=69-2=67 for the other angle.
Proof:
23+67=90
90=90

Answer by Earlsdon(6294)

The first angle is x degrees and the complementary angle is (3x-2), so...
x + (3x-2) = 90 Simplify.
4x - 2 = 90 Add 2 to both sides.
4x = 92 Divide both sides by 4
x = 23 degrees.
3x-2 = 3(23)-2 = 69-2 = 67 degrees.
Check:
23 + 67 = 90 degrees.